Slow Growth Company just paid an annual dividend of $1.65 a share. The firm expects to pay dividends forever and to increase the dividend by 3 percent annually. What is the expected value of this stock five years from now if the discount rate is 14%? Is it A$17.23, B.$19.88, C.$18.25, D.$17.91?
Answer #1

Present value = D1 / required rate - growth rate

Present value = [1.65(1 + 3%)] / 0.14 - 0.03

Present value = 1.6995 / 0.11

Present value = 15.45

Expected stock value in 5 years = Present value(1 + g)n

Expected stock value in 5 years = 15.45 (1 + 0.03)5

Expected stock value in 5 years = 15.45 * 1.159274

Expected stock value in 5 years = $17.91
Answer #2

present value = 1.65(1.03)/0.14-0.03

=15.45

expected price= 15.45(1.03)^5= 17.91


answered by: Ayyas